# A Concurrent Generalized Kropina Change

### v1

@@
Abstract. This paper investigates a generalized Kropina metric featuring a specific $\pi$-form. Start with a Finsler manifold $(M,F)$ admits a concurrent $\pi$-vector field $\overline{\varphi}$, then, examine the $\phi$-concurrent generalized Kropina change defined by $\widehat{F}=\frac{F^{m+1}}{\Phi^{m}}$, where $\Phi$ represents the corresponding $1$-form. We investigate the fundamental geometric objects associated with $\widehat{F}$ in an intrinsic manner after adopting this modification and present an example of a Finsler metric that admits a concurrent vector field along with $\widehat{F}$. Also, we prove that the geodesic sprays of $F$ and $\widehat{F}$ can never be projectively related. Moreover, we show $\overline{\varphi}$ is not concurrent with respect to $\widehat{F}$. Eventhough, we give a sufficient condition for $\overline{\varphi}$ to be concurrent with respect to $\widehat{F}$. Finally, we prove that the $\phi$-concurrent generalized Kropina change ($F\longrightarrow\widehat{F}$) preserves the almost rational property of the initial Finsler metric ${F}$.
@@ Introduction
Special Finsler spaces investigated globally in [5, 6, 7, 8, 17] and locally in [3, 11, 12]. Both local and global point of views are useful in treating problems in Finsler geometry. In this paper, we examine a Finsler manifold $(M,F)$ that possesses a concurrent $\pi$-vector field $\overline{\varphi}$. We use its corresponding $\pi$-form $\phi:=i_{\overline{\varphi}}g$, where $g$ represents the metric tensor of $F$, leading to the associated function $\Phi(x,y):=\phi(\overline{\eta})$. Next, we examine what we called the $\phi$-concurrent generalized Kropina change $\widehat{F}=F^{m+1}\Phi^{-m}$. In this context, we compute intrinsic geometric objects related to $\widehat{F}$. Namely, the supporting form $\widehat{\ell}$, the angular metric tensor $\widehat{\hbar}$, the Finsler metric $\widehat{g}$, the Cartan torsion $\widehat{\mathbf{T}}$, the geodesic spray $\widehat{G}$, the nonlinear connection $\widehat{\Gamma}$, the Berwald connection and the curvature tensor $\widehat{\Re}$ associated with $\widehat{\Gamma}$ are identified in terms of the corresponding geometric objects of $F$. Furthermore, we characterise the non-degenerate property of the metric tensor $\widehat{g}$ in §2.
We have noted that the above mentioned geometric objects are not invariant under the $\phi$-concurrent generalized Kropina change except the vertical counterpart of Berwald connection. We find a sufficient condition that makes some geometric objects, namely, $\widehat{\Gamma},\,\Re$ and the horizontal counterpart of Berwald connection to be invariant. Consequently, we prove that the $\pi$-vector field $\overline{\varphi}$ is not concurrent with respect to $\widehat{F}$, however, it will be under certain condition. On the other hand, we conclude the geodesic sprays $G$ and $\widehat{G}$ can never be projectively related. An example which represent our change is provided. We end this work by study the effect of the $\phi$-concurrent generalized Kropina change on an almost rational Finsler metric and obtain interesting results.
@@ 1 Preliminaries
$0\longrightarrow\pi^{-1}(TM)\stackrel{{\scriptstyle\gamma}}{{\longrightarrow}}TTM\stackrel{{\scriptstyle\rho}}{{\longrightarrow}}\pi^{-1}(TM)\longrightarrow 0,\vspace{-0.1cm}$
@@
For a linear connection $D$ on $\pi^{-1}(TM)$, we have $K:TTM\longrightarrow\pi^{-1}(TM)$ which is defined by $K(W)=D_{W}\overline{\eta}$. Thereby, the horizontal space at $u\in TM$ is $H_{u}(TM):=\{W\in T_{u}(TM)\,|\,K(W)=0\}$. The connection $D$ is said to be regular if for all $u\in TM$, we have $T_{u}(TM)=V_{u}(TM)\oplus H_{u}(TM).$ For a regular connection $D$, the vector bundle maps $\rho|_{H(TM)}$ and $K|_{V(TM)}$ are isomorphisms. In this case, the map $\beta:={\rho^{-1}}|_{H(TM)}$ is called the horizontal map of $D$. A well-Known regular connection is Berwald connection [2] which can be defined by [16, Proposition 4.4]
@@
Moreover, a spray on $M$ is a smooth vector field $G$ on ${\cal T}M$ such that $JG=\mathcal{C}$ and $[\mathcal{C},G]=G$. It is clear that $G=\beta\,\overline{\eta}$. A nonlinear connection on $M$ is a vector $1$-form $\Gamma$ on $TM$ which is smooth on ${\cal T}M$ and continuous on $TM$ such that $J\Gamma=J\text{ and }\,\,\Gamma J=-J$ [2]. Consequently, the horizontal and vertical projectors associated with $\Gamma$ are given, respectively, by
@@
Consequently, we get $vJ=J$ and $Jv=0$. The curvature of $\Gamma$ is defined by $\mathfrak{R}:=-\frac{1}{2}[h,h]$, which can be computed using Frölicher-Nijenhuis bracket $[\mathbb{K},\mathbb{L}]$ of two vector 1-forms $\mathbb{K}$ and $\mathbb{L}$ as follows [1]:
@@ A Finsler metric on $M$ is a function $F:TM\longrightarrow[0,\infty)$ such that:
The Finsler metric tensor $g$ induced by $\,F\,$ on $\pi^{-1}(TM)$ is defined as follows [18]
@@
In this case, the pair $(M,F)$ is called a Finsler manifold and $F$ is a regular Finsler metric. If the above conditions are satisfied on a conic subset of $TM$, then $(M,F)$ is called a conic Finsler manifold.
@@ 2 $\phi$-concurrent generalized Kropina change
$\bullet$ $||\overline{\varphi}||_{g}:=g(\overline{\varphi},\overline{\varphi})=\phi(\overline{\varphi})$ is the length of $\overline{\varphi}$ with respect to $F$.
Remark 2.2.
A concurrent $\pi$-vector field $\overline{\varphi}$ on a Finsler manifold $(M,F)$ and its corresponding $\pi$-form $\phi$ have no dependence of the directional argument $y$ [18, Theorem 3.7]. That is,
@@
$i_{\gamma\overline{\varphi}}\,J=0,\qquad d_{J}||\overline{\varphi}||_{g}=0,\qquad[\gamma\overline{\varphi},J]=0.$ (2.3)
@@
with $m\neq 0,-1$. The change ($F\longrightarrow\widehat{F}$) is called the $\phi$-concurrent generalized Kropina change. If $\widehat{F}$ is a Finsler metric on $M$, we are going to be called a $\phi$-concurrent generalized Kropina metric.
@@
The supporting form $\widehat{\ell}$ and $\ell$ are related by
^$\widehat{\ell}(\overline{X})={F}^{m}{\Phi}^{-m}\left\{(m+1)\ell(\overline{X})-m{F}{\Phi}^{-1}\phi(\overline{X})\right\}.$ (2.8)
@@
^$\widehat{\hbar}(\overline{X},\overline{Y})$ $=$ $(m+1){F}^{2m}{\Phi}^{-2m}\big{\{}(\hbar(\overline{X},\overline{Y})+m\ell(\overline{X})\,\ell(\overline{Y}))$
@@
(1) As the horizontal map ${\widehat{\beta}}$ of $\widehat{F}$ can be written in terms of the the horizontal map $\beta$ of $F$ (in the form $\widehat{\beta}=\beta+\gamma\overline{U}$, for some $\overline{U}$) and we have $\rho\circ\gamma=0$ in addition to ${{{D}}}^{\circ}_{\gamma\overline{X}}\overline{Y}=\rho[\gamma\overline{X},{\beta}\overline{Y}]$, from (1.3), therefore, we obtain
@@
[…]
@@
^$\widehat{\ell}(\overline{X})$ $=$ $d_{J}\widehat{F}(\widehat{\beta}\overline{X})=d_{J}\widehat{F}({\beta}\overline{X})=\frac{\partial\widehat{F}}{\partial F}\,d_{J}F(\beta\overline{X})+\frac{\partial\widehat{F}}{\partial\Phi}\,d_{J}\Phi(\beta\overline{X})$
@@
^$\widehat{\hbar}(\overline{X},\overline{Y})$ $=$ $\widehat{F}({{\widehat{D}}}^{\circ}_{\gamma\overline{X}}\,\widehat{\ell})(\overline{Y})=\widehat{F}({{D}}^{\circ}_{\gamma\overline{X}}\,\widehat{\ell})(\overline{Y})$
@@
$=$ $\widehat{F}\,\left\{({{D}}^{\circ}_{\gamma\overline{X}}\,\left((m+1)\Phi^{-m}F^{m}\right))\,\ell(\overline{Y})-({{D}}^{\circ}_{\gamma\overline{X}}\,m\Phi^{-m-1}F^{m+1})\,\phi(\overline{Y})\right\}$
$+\widehat{F}\,\left\{(m+1)\Phi^{-m}F^{m}\,({{D}}^{\circ}_{\gamma\overline{X}}\,\ell)(\overline{Y})-m\Phi^{-m-1}F^{m+1}({{D}}^{\circ}_{\gamma\overline{X}}\,\phi)(\overline{Y})\right\}$
@@
(5) Using items (3), (4) above and the definition ^$\widehat{\hbar}:=\widehat{g}-\widehat{\ell}\otimes\widehat{\ell}$, we obtain
@@
+$+\left\{(m+1)\Phi^{-m}F^{m}\,\ell(\overline{X})-m\Phi^{-m-1}F^{m+1}\,\phi(\overline{X})\right\}\,\times$
@@
$mF^{2}||\overline{\varphi}||_{g}-(m-1)\Phi^{2}\neq 0.$ (2.10)
In other words, the Finsler metric tensor $\widehat{g}$ of $\widehat{F}$ is non-degenerate if and only if the function $(mF^{2}||\overline{\varphi}||_{g}-(m-1)\Phi^{2})$ does not vanish identically.
@@
00 0 $=$ $(m+1)\Phi^{-2m}F^{2m}g(\overline{U},\overline{V})+m(2m+1)\Phi^{-2(m+1)}F^{2m+2}\,\phi(\overline{U})\,\phi(\overline{V})$
@@
where $\mathcal{A}_{1}:=2m(m+1)\Phi^{-2m-1}F^{2m-1}(\Phi^{2}-F^{2}||\overline{\varphi}||_{g})$,
$\mathcal{B}_{1}$ ::=:= $\Phi^{-2(m+1)}F^{2m}\left(F^{2}m(2m+1)||\overline{\varphi}||_{g}-\Phi^{2}\left(2m^{2}+m-1\right)\right),$
$\mathcal{A}_{2}$ ::=:= $(m+1)\Phi^{-2m}F^{2m+1},\,\qquad\mathcal{B}_{2}:=-m\,\Phi^{-2m-1}F^{2m+2}.$
@@
$\overline{Y}\neq 0\quad\Longleftrightarrow\quad mF^{2}||\overline{\varphi}||_{g}-(m-1)\Phi^{2}=0.$
@@
Form now on, we consider that the $\phi$-generalized Kropina metric $\widehat{F}$ satisfies the condition (2.10). Let us give an example of the $\phi$-generalized Kropina change of a Finsler metric which admits a $\pi$-concurrent vector field. Its calculations can be done by hand or using the Finsler package [20].
@@
Let $M=\{x=(x^{1},x^{2},x^{3})\in\mathbb{R}^{3}:\,x^{1}>0\}$ and consider the conic Finsler metric defined on the domain $\mathcal{D}=\{(x,y)\in TM|y^{3}\neq 0\}$ by
@@
Clearly, the non-vanishing components $g_{ij}$ of the metric tensor are
@@
Now, consider the vector field given by the components
@@
which defines a $\phi$-concurrent generalized Kropina metric over $M$.
@@
where $\Psi_{1}:=\frac{2m\Phi F^{2}}{mF^{2}||\overline{\varphi}||_{g}-(m-1)\Phi^{2}}$, $\Psi_{2}:=\frac{mF^{4}}{mF^{2}||\overline{\varphi}||_{g}-(m-1)\Phi^{2}}$.
@@
$-d\widehat{E}=i_{\widehat{G}}dd_{J}\widehat{E}.$
@@
00 0 $=$ $(m+1)[\Phi F^{-1}g(\overline{\mu},\rho X)+2m\ell(\rho{X})(F^{2}+\Phi^{-1}F\ell(\overline{\mu})-\phi(\overline{\mu}))-2m\phi(\rho{X})\ell(\overline{\mu})]$
@@
$A_{1}\,\ell(\overline{\mu})+B_{1}\,\phi(\overline{\mu})$ $=$ $\,C_{1},\qquad A_{2}\,\ell(\overline{\mu})+B_{2}\>\phi(\overline{\mu})=\,C_{2},$ (2.21)
@@
$A_{1}$ ::=:= $(m+1),\,\,\quad B_{1}:=-m\Phi^{-1}F,\,\,\quad C_{1}:=-m\Phi^{-1}F^{3},$
$A_{2}$ ::=:= $2m(m+1)\Phi^{-1}F^{-1}\left(\Phi^{2}-F^{2}||\overline{\varphi}||_{g}\right),$
$B_{2}$ ::=:= $\Phi^{-2}\left(F^{2}m(2m+1)||\overline{\varphi}||_{g}-\Phi^{2}\left(2m^{2}+m-1\right)\right),$
$C_{2}$ ::=:= $m\Phi^{-2}F^{2}\left(F^{2}(2m+1)||\overline{\varphi}||_{g}-2\Phi^{2}(m+1)\right).$
@@
$\ell(\overline{\mu})$ $=$ $\frac{m\Phi F^{3}}{(m-1)\Phi^{2}-mF^{2}||\overline{\varphi}||_{g}},\quad\phi(\overline{\mu})=\frac{mF^{2}\left(F^{2}||\overline{\varphi}||_{g}-2\Phi^{2}\right)}{mF^{2}||\overline{\varphi}||_{g}-(m-1)\Phi^{2}}.$
@@
$\widehat{G}=G-\frac{2m\Phi F^{2}}{mF^{2}||\overline{\varphi}||_{g}-(m-1)\Phi^{2}}\,{\mathcal{C}}+\frac{mF^{4}}{mF^{2}||\overline{\varphi}||_{g}-(m-1)\Phi^{2}}\,\gamma\overline{\varphi}.$
@@
The Barthel curvature tensors $\widehat{\Re}$ and $\Re$ are determined by ^$\widehat{\Re}=\Re-\frac{1}{2}[h,\mathbb{F}]-\frac{1}{4}N_{\mathbb{F}}.$
@@
The Barthel curvature tensor of $\widehat{F}$ is defined by ^$\widehat{\Re}=-\frac{1}{2}[\widehat{h},\widehat{h}]$. Now, item (1) above, (2.23) and formula (1.6) together with the aspects of the Frölicher-Nijenhuis bracket lead to
@@
${=}$ $\gamma{D^{\circ}}_{h{W}}\overline{Z}+\frac{\gamma}{2}\big{\{}-\Psi_{1}\,K[\,JW,JZ]-d_{J}\Psi_{1}(W)\,K[\,\gamma\overline{\eta},JZ]$

### v2

@@
Abstract. This paper investigates a generalized Kropina metric featuring a specific $\pi$-form. Start with a Finsler manifold $(M,F)$ admits a concurrent $\pi$-vector field $\overline{\varphi}$, then, examine the $\phi$-concurrent generalized Kropina change defined by $\widehat{F}=\frac{F^{m+1}}{\Phi^{m}}$, where $\Phi$ represents the corresponding $1$-form, $\Phi^{m}>0$. We investigate the fundamental geometric objects associated with $\widehat{F}$ in an intrinsic manner after adopting this modification and present an example of a Finsler metric that admits a concurrent vector field along with $\widehat{F}$. Also, we prove that the geodesic sprays of $F$ and $\widehat{F}$ can never be projectively related. Moreover, we show $\overline{\varphi}$ is not concurrent with respect to $\widehat{F}$. Eventhough, we give a sufficient condition for $\overline{\varphi}$ to be concurrent with respect to $\widehat{F}$. Finally, we prove that the $\phi$-concurrent generalized Kropina change ($F\longrightarrow\widehat{F}$) preserves the almost rational property of the initial Finsler metric F{F}.
@@ Introduction
Special Finsler spaces investigated globally in [5, 6, 7, 8, 18] and locally in [3, 11, 12, 13]. Both local and global point of views are useful in treating problems in Finsler geometry. In this paper, we examine a Finsler manifold $(M,F)$ that possesses a concurrent $\pi$-vector field $\overline{\varphi}$. We use its corresponding $\pi$-form $\phi:=i_{\overline{\varphi}}g$, where $g$ represents the metric tensor of $F$, leading to the associated function $\Phi(x,y):=\phi(\overline{\eta})$. Next, we examine what we called the $\phi$-concurrent generalized Kropina change $\widehat{F}=F^{m+1}\Phi^{-m}$. In this context, we compute intrinsic geometric objects related to $\widehat{F}$. Namely, the supporting form $\widehat{\ell}$, the angular metric tensor $\widehat{\hbar}$, the Finsler metric $\widehat{g}$, the Cartan torsion $\widehat{\mathbf{T}}$, the geodesic spray $\widehat{G}$, the nonlinear connection $\widehat{\Gamma}$, the Berwald connection and the curvature tensor $\widehat{\Re}$ associated with $\widehat{\Gamma}$ are identified in terms of the corresponding geometric objects of $F$. Furthermore, we characterise the non-degenerate property of the metric tensor $\widehat{g}$ in §2.
We have noted that the above mentioned geometric objects are not invariant under the $\phi$-concurrent generalized Kropina change except the vertical counterpart of Berwald connection. We find a sufficient condition that makes some geometric objects, namely, $\widehat{\Gamma}$$\Re$ and the horizontal counterpart of Berwald connection to be invariant. Consequently, we prove that the $\pi$-vector field $\overline{\varphi}$ is not concurrent with respect to $\widehat{F}$, however, it will be under certain condition. On the other hand, we conclude the geodesic sprays $G$ and $\widehat{G}$ can never be projectively related. An example which represent our change is provided. We end this work by study the effect of the $\phi$-concurrent generalized Kropina change on an almost rational Finsler metric and obtain interesting results.
@@ 1 Preliminaries
$0\longrightarrow\pi^{-1}(TM)\stackrel{{\scriptstyle\gamma}}{{\longrightarrow}}TTM\stackrel{{\scriptstyle\rho}}{{\longrightarrow}}\pi^{-1}(TM)\longrightarrow 0,\vskip-2.84544pt$
@@
For a linear connection $D$ on $\pi^{-1}(TM)$, we have $K:TTM\longrightarrow\pi^{-1}(TM)$ which is defined by $K(W)=D_{W}\overline{\eta}$. Thereby, the horizontal space at $u\in TM$ is $H_{u}(TM):=\{W\in T_{u}(TM)\,|\,K(W)=0\}$. The connection $D$ is said to be regular if for all $u\in TM$, we have $T_{u}(TM)=V_{u}(TM)\oplus H_{u}(TM).$ For a regular connection $D$, the vector bundle maps $\rho|_{H(TM)}$ and $K|_{V(TM)}$ are isomorphisms. In this case, the map $\beta:={\rho^{-1}}|_{H(TM)}$ is called the horizontal map of $D$. A well-Known regular connection is Berwald connection [2] which can be defined by [17, Proposition 4.4]
@@
[…]
@@
Consequently, we get $vJ=J$ and $Jv=0$. The curvature of $\Gamma$ is defined by $\mathfrak{R}:=-\frac{1}{2}[h,h]$, which can be computed using Frölicher-Nijenhuis bracket $\mathbb{K},\mathbb{L}]$ of two vector 1-forms $\mathbb{K}$ and $\mathbb{L}$ as follows [1]:
@@ A Finsler metric on $M$ is a function $F:TM\longrightarrow[0,\infty)$ such that:
The Finsler metric tensor $g$ induced by F\,F\, on $\pi^{-1}(TM)$ is defined as follows [19]
@@
In this case, the pair $(M,F)$ is called a Finsler manifold and $F$ is a Finsler metric.
If the above conditions are satisfied on a conic subbundle $\mathbf{D}$ of $TM$, then $(M,F)$ is called a conic pseudo-Finsler manifold. A conic subbundle of $TM$ is a non-empty open subset $\mathbf{D}\subset{\cal T}M$ that is invariant under the scaling of its tangent vectors by positive real numbers. Additionally, it must satisfy the condition $\pi(\mathbf{D})=M$. Furthermore, it is assumed that for every $x\in M$, the set $\mathbf{D}_{x}:=\mathbf{D}\cap T_{x}M$ is a connected set.
@@ 2 $\phi$-concurrent generalized Kropina change
$\bullet$ $\overline{\varphi}||_{g}:=g(\overline{\varphi},\overline{\varphi})=\phi(\overline{\varphi})$ is the length of $\overline{\varphi}$ with respect to $F$.
Lemma 2.3.
A concurrent $\pi$-vector field $\overline{\varphi}$ on a Finsler manifold $(M,F)$ and its corresponding $\pi$-form $\phi$ have no dependence of the directional argument $y$ [19, Theorem 3.7]. That is,
@@
[…]
@@
with $m\neq 0,-1$. The change ($F\longrightarrow\widehat{F}$) is called the $\phi$-concurrent generalized Kropina change. If $\widehat{F}$ is a conic pseudo-Finsler metric on $M$, then it is going to be called a $\phi$-concurrent generalized Kropina metric.
@@
The supporting elements $\widehat{\ell}$ and $\ell$ are related by
$\widehat{\ell}(\overline{X})={F}^{m}{\Phi}^{-m}\left\{(m+1)\ell(\overline{X})-m{F}{\Phi}^{-1}\phi(\overline{X})\right\}.$ (2.8)
@@
$\widehat{\hbar}(\overline{X},\overline{Y})$ $=$ $(m+1){F}^{2m}{\Phi}^{-2m}\big\{(\hbar(\overline{X},\overline{Y})+m\ell(\overline{X})\,\ell(\overline{Y}))$
@@
(1) As the horizontal map $^{\widehat{\beta}}$ of $\widehat{F}$ can be written in terms of the the horizontal map $\beta$ of $F$ (in the form $\widehat{\beta}=\beta+\gamma\overline{U}$, for some $\overline{U}$) and we have $\rho\circ\gamma=0$ in addition to ${{{D}}}^{\circ}_{\gamma\overline{X}}\overline{Y}=\rho[\gamma\overline{X},{\beta}\overline{Y}]$, from (1.3), therefore, we obtain
@@
$\textbf{ (2)}\,Clearly,\,\,d\widehat{E}(Z)$ $=$ $\frac{1}{2}d\widehat{F}^{2}(Z)=\widehat{F}\,d\widehat{F}(Z)$
@@
$\widehat{\ell}(\overline{X})$ $=$ $d_{J}\widehat{F}(\widehat{\beta}\overline{X})=d_{J}\widehat{F}({\beta}\overline{X})=\frac{\partial\widehat{F}}{\partial F}\,d_{J}F(\beta\overline{X})+\frac{\partial\widehat{F}}{\partial\Phi}\,d_{J}\Phi(\beta\overline{X})$
@@
$\widehat{\hbar}(\overline{X},\overline{Y})$ $=$ $\widehat{F}({{\widehat{D}}}^{\circ}_{\gamma\overline{X}}\,\widehat{\ell})(\overline{Y})=\widehat{F}({{D}}^{\circ}_{\gamma\overline{X}}\,\widehat{\ell})(\overline{Y})$
@@
[…]
@@
(5) Using items (3), (4) above and the definition $\widehat{\hbar}:=\widehat{g}-\widehat{\ell}\otimes\widehat{\ell}$, we obtain
@@
$+\left\{(m+1)\Phi^{-m}F^{m}\,\ell(\overline{X})-m\Phi^{-m-1}F^{m+1}\,\phi(\overline{X})\right\}\,\times$
@@
[…]
@@
$0$ $=$ $(m+1)\Phi^{-2m}F^{2m}g(\overline{U},\overline{V})+m(2m+1)\Phi^{-2(m+1)}F^{2m+2}\,\phi(\overline{U})\,\phi(\overline{V})$
@@
[…]
$\mathcal{A}_{2}$ $:=$ $(m+1)\Phi^{-2m}F^{2m+1},\,\qquad\mathcal{B}_{2}:=-m\,\Phi^{-2m-1}F^{2m+2}.$
@@
[…]
@@
Form now on, we consider that the $\phi$-generalized Kropina metric $\widehat{F}$ satisfies the condition (2.10) and $\Phi^{m}>0$ in order to make $\widehat{F}$ Finsler metric. In fact, $\widehat{F}$ is a conic pseudo-Finsler metric. Let us give an example of the $\phi$-generalized Kropina change of a Finsler metric which admits a $\pi$-concurrent vector field. Its calculations can be done by hand or using the Maple Finsler package [21].
@@
Let $M=\{x:=(x^{1},x^{2},x^{3})\in\mathbb{R}^{3}:\,x^{1}>0\}$ and consider the conic Finsler metric defined on the domain $\mathcal{D}=\{(x,y:=y^{1},y^{2},y^{3}))\in TM|\,y^{3}\neq 0\}$ by
@@
Clearly, the non-vanishing components $g_{ij}$ of the Finsler metric tensor are
@@
Now, consider the vector field whose components given by
@@
which defined on the domain:
$\mathbf{D}=\mathcal{D}\cap\{(x,y)\in TM|\,(y^{1})^{m}>0\}.$
Assume that $\widehat{F}(x,y)$ satisfies the condition (2.10), hence it is a $\phi$-concurrent generalized Kropina metric.
@@
[…]
@@
$d\widehat{E}=i_{\widehat{G}}dd_{J}\widehat{E}$.
@@
$0$ $=$ $(m+1)[\Phi F^{-1}g(\overline{\mu},\rho X)+2m\ell(\rho{X})(F^{2}+\Phi^{-1}F\ell(\overline{\mu})-\phi(\overline{\mu}))-2m\phi(\rho{X})\ell(\overline{\mu})]$
@@
$A_{1}\,\ell(\overline{\mu})+B_{1}\,\phi(\overline{\mu})$ $=$ $\,C_{1},\hskip 18.49988ptA_{2}\,\ell(\overline{\mu})+B_{2}\>\phi(\overline{\mu})=\,C_{2},$ (2.21)
@@
$A_{1}$ $:=$ $(m+1),\,\,\qquad B_{1}:=-m\Phi^{-1}F,\,\,\qquad C_{1}:=-m\Phi^{-1}F^{3},$
[…]
@@
[…]
@@
[…]
@@
The Barthel curvature tensors $\widehat{\Re}$ and $\Re$ are determined by $\widehat{\Re}=\Re-\frac{1}{2}[h,\mathbb{F}]-\frac{1}{4}N_{\mathbb{F}}.$
@@
The Barthel curvature tensor of $\widehat{F}$ is defined by $\widehat{\Re}=-\frac{1}{2}[\widehat{h},\widehat{h}$]. Now, item (1) above, (2.23) and formula (1.6) together with the aspects of the Frölicher-Nijenhuis bracket lead to
@@
$\gamma\widehat{D^{\circ}}_{h{W}}\,\rho{Z}$ $=$ $\widehat{v}\,[\widehat{h}W,JZ]\overset{\textbf{(2)}}{=}(v-\frac{1}{2}\mathbb{L})[(h+\frac{1}{2}\mathbb{L})W,JZ]$